The coin is in shape of a circle of radius 3cm with a square of sides x cm removed from its centre. The area of each face of the coin in 7πcm^2.
i) form an equation in x and show that it reduces to 2π-x^2=0.
ii) solve the equation 2π-x^2=0.
iii) find the perimeter of the square.

Answer by Theo(13342)

this was hard to visualize, but i finally got it right, i think.

algebraically is's not that difficult.

what you have is a circle with a radius of 3 centimeters.

since the area of a circle is equal to pi * r^2, then the area of this circle is 9 * pi.

you are told that the area of the circle minus the area of the square is equal to 7 * pi.

you are told that the length of a side of the square is x.

that means the area of the square is x^2.

therefore, the area of the circle minus the area of the square must be equal to 9 * pi - x^2.

if the net area is 7 * pi, then you get:

9 * pi - x^2 = 7 * pi

subtract 7 * pi from both sides of this equation and you get:

9 * pi - 7 * pi - x^2 = 0

this reduces to 2 * pi - x^2 = 0

this is the formula you are looking for.

if you solve this formula for x^2, you will x^2 = 2 * pi.

the area of the square inside the circle is 2 * pi.

that makes each side of the square equal to sqrt(2 * pi).

the perimeter of the square is 4 * the length of each side of the square.

that makes the perimeter of the square equal to 4 * sqrt(2 * pi).

you can play games with this answer, but it will be the same.

for example:

4 = sqrt(16), therefore the answer becomes sqrt(16) * sqrt(2 * pi) which then becomes sqrt(16 * 2 * pi) which becomes sqrt(32 * pi).

if i did this right, 4 * sqrt(2 *pi) should be equal to sqrt(32 * pi).

i used my calculator to confirm that they are, indeed, equal to each other.

so, your answer can be that the perimeter of the square is either:

4 * sqrt(2 * pi) or:
sqrt(32 * pi) or, if you need a decimal answer:
10.0265131 centimeters.

my problem was in visualizing the answer.
i finally got that right.
a graphical display of the circular coin and the square in the center of the coin would look like this:

$$$

you can see from the graph, that the length of each side of the square is 2 * 1.25331413732 = 2.506628275.

square this to get 2.506628275^2 = 6.283185307.

multiply this by pi / pi and it becomes 6.283185307 * (pi / pi) which is the same as (6.283185307 / pi) * pi which can be reduced to 2 * pi.

in other words, 6.283185307 is equivalent to 2 * pi, which is the area of the square that we had calculated to be earlier.